You are bullish on Curious Inc. stock and decide to purchase 200 shares at $30 per share. To do the transaction, you borrowed $2,000 from your broker on 9% interest. What is the remaining margin on your account, if the share price drops to $17.3 a year later?(Provide your answer in % rounded to two digits omitting the % sign)
Answer #1

Initial margin = 200 * 30 - 2000 = 4000

Decrease in portfolio value = (30 -17.3) *20 = 2540

Remaining balance = 4000 - 2540 = 1460

In percentage terms = 1460 / 4000 = 36.5%